OpenStudy (jmansedg):

is 5 a rational number a true b false

OpenStudy (anonymous):

Rational numbers can be written as fractions. Can 5 be written as a fraction?

OpenStudy (jmansedg):

yes 2/3

OpenStudy (anonymous):

you're right with the yes, by not the 2/3. 5 = 5/1 = 10/2 = 15/3, etc

OpenStudy (jmansedg):

ok thanks

OpenStudy (anonymous):

yw